Ponn Inc. needs to borrow $250,000 for the next 6 months. The company has a line of credit with a bank that allows the company to borrow funds with an compensating balance. Currently, Penn Inc. has no funds on deposit with the bank and will need the loan to cover the compensating balance as well as their other financing needs. What will be the interest rate subject to a 20% of loan total interest amount for this financing? 

  • $11,250 

  • $10,000 

  • $12.500 

  • $8,750
Answer #1

amount of loan required for 6months = 250000

interest rate = 7% per anum provided 20% of the loan shall be the compensated balance

20% * 250000 = 50000

penn inc. doesnt have any funds to compensate . therefore the same shall be borrowed

the amount of 50000 shall be borrowed at a rate higher than the 7%

therefor option B & D are not correct as explained below

explaination : if 50000 borrowed @ 7% the the toal interest payable = (250000+50000)*7%*6/12= 10500 but the interest rate o 50000 shall be more than 7%. the interest payable shall be more than 10500 but not less than such amount. therefore option B &D not correct.

due to lack of information in question it is assumed interest rate on 50000 be 10%.

the total interest payable = [250000*7%*6/12] + [50000*10%*6/12] = 11250